Incorrect Question 5 0/1 pts According to NASA, an average of one cataloged piece of space debris has fallen back to Earth each day for the past 50 years. Suppose that there is an 11.5% chance that a piece of debris will fall in Russia during a given week, and a 6.7% chance that a piece of debris will fall in Canada during the same week. Which is the most likely? Debris will fall only in Russia Debris will fall in Russia and in Canada Debris will fall only in Canada Debris will fall in Russia or in Canada.
Answer #1

Let R shows the event that Debris will fall in Russia and C shows the event that Debris will fall in Canada.

So we have

P(R) = 0.115, P(C) = 0.067

Since these events are independent so

P(R and C) = P(R)P(C) = 0.115 * 0.067 = 0.007705

--------------

P(only R) = P(R) - P(R and C) = 0.115 - 0.007705 = 0.107295

------------------

P(only C) = P(C) - P(R and C) = 0.067- 0.007705 = 0.059295

--------------

P(R or C) = P(R) + P(C) - P(R and C) = 0.115 + 0.067 - 0.007705 = 0.174295

Hence, last option is correct.